Problem 17-55 (Algo) Derive Amounts for Profit Variance Analysis (LO 17-6)

Classics, Ltd., details cars. Classics wants to compare this quarter’s results with those for last quarter, which is believed to be typical for operations. Assume that the following information is provided.
 


Last Quarter
This Quarter
Number of detailings
440


523
Revenues$72,160

$68,800
Variable costs
27,920


31,920
Contribution margin$44,240

$36,880

 
Required:

a. Compute the flexible budget and sales activity variance and prepare a profit variance analysis. (Hint: Use last quarter as the master budget and this quarter as “actual.”)

b. What impact did the changes in number of detailings and average revenues (i.e., sales price) have on Classics, Ltd.’s contribution margin?
Answer #1

actual budget:

sales revenue 68,800 given

variable cost: 31,920 given

CM= 68,800- 31,920= 26,880


master budget

sales revenue: 72,160 given

variable cost: 27,920 given

CM= 72,160- 27,920= 44,240


Flexible budget:

sales revenue: 523* 72,260/ 440= 85,772

variable cost: 523* 27,920/ 440= 33,186. 727272727 or 33,187

contribution marigin (CM)= 85,772- 33,187= 52,585


variable cost variance: 

variable cost and CM= 33,187- 31,920= 1,267 F


sales price variance: 

sales revenue and CM= 68,800- 85,772= 16,972 U 

sales activity variance

for sales revenue, variable cost, and contribution margin

subtract flex budget from master budget


Part B) 16,972- 8,345- 1,267= 7,360 decrease by